OpenStudy (anonymous):

Solve for the variable. Reduce the answer to lowest terms. (7/6)s =91

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

s=1/78

OpenStudy (anonymous):

oops! its s=78

OpenStudy (anonymous):

Multipyling both sides by 6/7 (6/7) * (7/6)s = (6/7)*91 s = 546/7 = 78